Ship reports banner with links Home IODP-USIO daily/weekly reports

IODP Expedition 303:
North Atlantic Climate 1

Week 2 Report

PDF file is available for download.


October 4, 2004

Operations

Portcall: The port call concluded in St. John's, Newfoundland when the last line was released from the fueling dock (Pier 24) at 1925 hr on 30 September 2004. The departure was delayed because of two unassociated electrical problems related to the starboard shaft. Transocean also completed the final phase of an underwater hull inspection of a Special Periodical Survey #5 for the American Bureau of Shipping, which included the opening and inspection of 46 underwater hull valves, wind-waterline thickness gauging, anode replacement, cleaning, and inspection. At the conclusion of the inspection, the divers were able to clean and polish approximately 90% of the propeller surface area before it was necessary to secure the exercise when available time expired. While at the fueling dock, the vessel loaded 1678.454 MT of marine gasoil.

Transit: We completed the 276 nm transit from St. John's to Site 1302 at an average of 8.8 knots. During the transit, the ship's clocks were advanced 30 minutes to Mid-Atlantic Time (-2 hours UTC).

Site 1302 (ORPH3A)

Hole 1302A; 50 9.9845'N, 45 38.2713'W)
We arrived at Site 1302 at 0330 hr on 2 October and spudded Hole 1302A at 1530 hr. Piston coring, utilizing nonmagnetic core barrels, advanced to 107.1 mbsf (Cores 1H-13H) where detrital gravel caused problems with penetration and hole stability. An attempt was made to clean out the hole and continue coring, but hole stability problems persisted. Hole 1302A was terminated when the bit cleared the seafloor at 1245 hr. Operations included core orientation starting with Core 4H and two deployments of the APC temperature tool at 76.1 mbsf (Core 9H) and 104.6 mbsf (Core 12H). The latter attempt was unsuccessful because the cutting shoe and associated electronics were damaged by hard debris. At Hole 1302A we cored 107.1 m recovering 91.7 m (85.6%); recovery may have been affected by heave as large as 3 m.

Hole 1302B (50 9.952'N, 45 38.2897'W)
After offsetting the vessel 30 m to the northwest of Hole 1302A, Hole 1302B was spudded at 1340 hr on 3 October 2004. Piston coring advanced to 38.2 mbsf by 1800 hr on 3 October with an average recovery of 102%.

Technical Support and HSE Activities

The Expedition 303 technical staff began loading freight and preparing the laboratories for sea on 27 September 2004. All freight was loaded by the end of the day on 28 September 2004, except for a few port call purchases that were delivered on 29 September. A local Canon technician serviced the two shipboard copiers. All trash was collected and offloaded before departure. Equipment and laboratory spaces were secured for sea.

During the short transit to the first site, the technical crew prepared the shipboard laboratories for the commencement of coring. The scientific staff was introduced to the laboratories and technical staff.

HSE: A fire and boat drill was conducted on 2 October to familiarize new personnel with the location of their lifeboat. The technical staff joined other non-assigned personnel at the lifeboats.



About the IODP JRSO | Expeditions | Data & core samples | Tools & laboratories | Publications | Education |
Newsroom | Meetings & port calls | Employment | Contact us | Search | Site map | People | Intranet | Home

Accesibility Statement | Plug-in download | Privacy statement | ODP

For comments or questions: webmaster@iodp.tamu.edu

Copyright IODP JRSO

Modified on Wednesday, 31-May-2006 11:50:59 CDT.